# Patch-Level Tokenization with CNN Encoders and Attention for Improved Transformer Time-Series Forecasting

### v1

@@ Abstract
Transformer-based models have shown strong performance in time-series forecasting by leveraging self-attention to model long-range temporal dependencies. However, their effectiveness depends critically on the quality and structure of input representations derived from raw multivariate time-series data. This work proposes a two-stage forecasting framework that explicitly separates local temporal representation learning from global dependency modelling. In the first stage, a convolutional neural network (CNN) operates on fixed-length temporal patches to extract short-range temporal dynamics and non-linear feature interactions, producing compact patch-level token embeddings. Token-level self-attention is subsequently applied during representation learning to refine these embeddings by enabling interactions across temporal patches. In the second stage, a Transformer encoder processes the resulting token sequence to model inter-patch temporal dependencies and generate per-patch forecasts. Experiments conducted on synthetic multivariate time-series data with controlled static and dynamic factors demonstrate that the proposed patch-based tokenization strategy achieves competitive forecasting performance compared to convolutional and patch-based Transformer baselines. The results highlight the importance of structured temporal representations and show that decoupling local temporal encoding from global attention-based modelling yields more effective and stable time-series forecasting.
@@ I Introduction
Transformer-based architectures [14] have recently emerged as a powerful alternative for time-series forecasting. By employing self-attention mechanisms, Transformers enable direct interactions between all elements of a sequence, removing the reliance on recursive state propagation. This allows more effective modelling of long-range temporal dependencies and improved parallelization during training. Several studies have reported promising results using Transformer models for time-series forecasting [9, 15]. Nevertheless, Transformer performance is highly sensitive to the structure and quality of the input representations provided to the attention mechanism.
@@
The proposed method is evaluated on synthetic multivariate time-series datasets designed to reflect key structural properties commonly observed in real-world temporal data, including non-linear dynamics, interactions between static and dynamic factors, and stochastic noise. This controlled experimental setting enables systematic evaluation of the proposed representation learning strategy without introducing domain-specific confounding effects. Performance is compared against convolutional and patch-based Transformer baselines, including Temporal Convolutional Networks (TCN) and PatchTST, and results demonstrate that patch-level temporal tokenization consistently achieves competitive forecasting accuracy on unseen data.
@@ II-A Proposed method
II-B CNN Patch Encoder + Attention
@@ III-A1 Preparation
where $g(\cdot)$ denotes a non-linear transformation involving linear and logarithmic terms, $h(t)$ denotes a trend component, $p(t)$ denotes a periodic component, $\rho$ controls the dependency strength, and $\epsilon_{t}^{(2)}\sim\mathcal{N}(0,\sigma_{2}^{2})$.
@@
where $\alpha_{1}=0.045$, $\alpha_{2}=0.38$, $\alpha_{3}=0.07$, and $\epsilon_{t}^{(y)}\sim\mathcal{N}(0,\sigma_{y}^{2})$.
@@ III-A2 Seeding
To ensure a fair and reproducible evaluation, two independent datasets were generated using random seeds 42 and 101 for training and testing, respectively. Both datasets follow identical data generation and preprocessing pipelines. All proposed and baseline models are trained exclusively on the training dataset and evaluated on the held-out test dataset.
@@ III-B Compared Models
Temporal Convolutional Networks [1] employ stacks of one-dimensional convolutional layers with increasing receptive fields to model temporal dependencies in sequential data. By using dilated convolutions, TCNs are able to capture long-range temporal patterns while maintaining a fully convolutional structure and avoiding recurrent connections. This design enables stable training and parallel computation, making TCNs a strong baseline for time-series forecasting tasks. In our experiments, the TCN operates directly on the raw multivariate input sequence and produces patch-aligned forecasts using a comparable training protocol.
@@
PatchTST [11] is a recent Transformer-based time-series forecasting model that introduces a patch-level tokenization strategy prior to attention modelling. Instead of applying self-attention directly to individual time steps, the input sequence is partitioned into fixed-length temporal patches, which are linearly embedded and treated as tokens for a Transformer encoder. This approach reduces sequence length, improves computational efficiency, and encourages the model to focus on patch-level temporal structure. PatchTST serves as a strong baseline for evaluating the effectiveness of patch-based representations in Transformer architectures.
@@ III-C Implementation details
The proposed model and other compared models have been implemented in python with the PyTorch library; architecture and experimentation is publicly available [10]. All models are trained using the AdamW optimizer.
@@ III-D Hyperparameter Settings
Number of patches ($\lfloor T/P\rfloor$) 3
@@
Tables II and III summarize the global and model-specific hyperparameter settings used in the experiments
@@ IV Results
Figure 2 compares the forecasting performance of the proposed model with the TCN and PatchTST baselines. All models achieve low error values, indicating that the synthetic dataset is well suited for evaluating patch-based forecasting approaches. Among the evaluated methods, PatchTST attains the lowest error values, achieving an MSE of 0.028 and an MAE of 0.131. The proposed model achieves an MSE of 0.093 and an MAE of 0.232, while the TCN baseline records an MSE of 0.038 and an MAE of 0.153.
@@
The observed differences in performance can be interpreted in light of the architectural characteristics of the models. PatchTST applies self-attention directly to patch-level representations that are learned end-to-end for the forecasting objective, which may be advantageous in the considered synthetic setting where target values depend primarily on contemporaneous dynamic features and static covariates. This direct optimization of patch embeddings enables efficient modelling of non-linear feature interactions.
The TCN baseline achieves competitive performance, highlighting the effectiveness of convolutional architectures with large receptive fields for capturing localized temporal structure and reducing noise. This suggests that a substantial portion of the predictive signal in the dataset is contained within short-range temporal patterns that can be effectively modelled without explicit attention mechanisms.
@@
The proposed two-stage framework demonstrates stable and consistent performance across both error metrics. By decoupling local temporal representation learning from global dependency modelling, the approach produces structured patch-level tokens that remain informative for downstream forecasting, even though the representation learning and forecasting stages are trained separately. While the proposed method does not outperform PatchTST on this benchmark, it achieves comparable performance to strong convolutional and Transformer-based baselines, supporting the effectiveness of patch-level temporal tokenization for multivariate time-series forecasting.
@@ V Conclusion
This study investigates a multi-stage forecasting architecture that explicitly separates patch-level temporal tokenization from global sequence modelling in multivariate time-series forecasting. By first learning compact patch-level token embeddings through a convolutional encoder and subsequently refining these representations using token-level self-attention prior to Transformer-based modelling, the proposed framework captures localized temporal structure while enabling global dependency modelling at a coarser temporal resolution.
Empirical results on a controlled synthetic dataset designed to reflect key structural properties of financial time series demonstrate that the proposed approach achieves competitive forecasting performance relative to strong convolutional and patch-based Transformer baselines. While a direct patch-based Transformer model achieves the lowest prediction errors on this benchmark, the proposed method performs comparably across both MSE and MAE, indicating that decoupling local temporal representation learning from global attention-based modelling does not degrade predictive accuracy.
These findings suggest that offloading local temporal representation learning from the Transformer may provide a structured and stable alternative to end-to-end attention-based modelling, particularly in settings where localized temporal patterns and feature interactions play an important role. The modular nature of the proposed framework also facilitates representation reuse and architectural flexibility, which may be advantageous in more complex or heterogeneous forecasting scenarios.
While the current implementation employs a convolutional patch encoder for token generation, future work may explore alternative patch-level encoders, including recurrent or hybrid architectures, as well as jointly trained or task-aware representation learning strategies. Although the evaluation is conducted on synthetic data, the results highlight the potential of structured patch-level tokenization for multivariate time-series forecasting, with the understanding that performance on real-world financial data may depend on domain-specific characteristics and market conditions.

### v2

@@ Abstract
Transformer-based models have shown strong performance in time-series forecasting by leveraging self-attention to model long-range temporal dependencies. However, their effectiveness depends critically on the quality and structure of input representations derived from raw multivariate time-series data, particularly as sequence length and data scale increase. This paper proposes a two-stage forecasting framework that explicitly separates local temporal representation learning from global dependency modelling. In the proposed approach, a convolutional neural network operates on fixed-length temporal patches to extract short-range temporal dynamics and non-linear feature interactions, producing compact patch-level token embeddings. Token-level self-attention is applied during representation learning to refine these embeddings, after which a Transformer encoder models inter-patch temporal dependencies to generate forecasts. The method is evaluated on a synthetic multivariate time-series dataset with controlled static and dynamic factors, using an extended sequence length and a larger number of samples. Experimental results demonstrate that the proposed framework consistently outperforms a convolutional baseline under increased temporal context and remains competitive with a strong patch-based Transformer model. These findings indicate that structured patch-level tokenization provides a scalable and effective representation for multivariate time-series forecasting, particularly when longer input sequences are considered.
@@ I Introduction
Transformer-based architectures [16] have recently emerged as a powerful alternative for time-series forecasting. By employing self-attention mechanisms [15], Transformers enable direct interactions between all elements of a sequence, removing the reliance on recursive state propagation. This allows more effective modelling of long-range temporal dependencies and improved parallelization during training. Several studies have reported promising results using Transformer models for time-series forecasting [9, 17]. Nevertheless, Transformer performance is highly sensitive to the structure and quality of the input representations provided to the attention mechanism.
@@
The proposed method is evaluated on synthetic multivariate time-series datasets designed to reflect key structural properties commonly observed in real-world temporal data, including non-linear dynamics, interactions between static and dynamic factors, and stochastic noise. This controlled experimental setting enables systematic evaluation of the proposed representation learning strategy without introducing domain-specific confounding effects. Performance is compared against convolutional and patch-based Transformer baselines, including Temporal Convolutional Networks (TCN) [1] and PatchTST [11], and results demonstrate that patch-level temporal tokenization consistently achieves competitive forecasting accuracy on unseen data.
The remainder of this paper is organized as follows. Section II describes the proposed patch-based forecasting framework, detailing the CNN-based patch encoder, token-level self-attention mechanism, and the specialized Transformer architecture used for global temporal modelling. Section III presents the experimental setup, including dataset generation, baseline models, implementation details, and hyperparameter settings. Section IV reports the empirical results and provides a comparative analysis of the proposed method against baseline approaches. Finally, Section V concludes the paper and discusses directions for future research.
@@ II-A Proposed method
II-B CNN Patch Encoder and Attention
@@ III-A1 Preparation
where $g(\cdot)$ denotes a fixed non-linear transformation composed of linear and logarithmic terms, $h(t)$ denotes a trend component, $p(t)$ denotes a periodic component, $\rho$ controls the dependency strength, and $\epsilon_{t}^{(2)}\sim\mathcal{N}(0,\sigma_{2}^{2})$.
@@
where $\alpha_{1}=0.045$, $\alpha_{2}=0.38$, $\alpha_{3}=0.07$, and $\epsilon_{t}^{(y)}\sim\mathcal{N}(0,\sigma_{y}^{2})$. The noise variances $\sigma_{1}^{2}$, $\sigma_{2}^{2}$, and $\sigma_{y}^{2}$ are fixed constants controlling the noise level in the synthetic data generation process.
@@ III-A2 Seeding
To ensure a fair and reproducible evaluation, two independent datasets were generated using random seeds 42 and 101 for training and testing, respectively. Both datasets follow identical data generation and preprocessing pipelines. There are 10,000 samples in each dataset. All proposed and baseline models are trained exclusively on the training dataset and evaluated on the held-out test dataset.
@@ III-B Compared Models
Temporal Convolutional Networks (TCNs) [1] model temporal dependencies using stacks of one-dimensional dilated convolutions, allowing large receptive fields without recurrent connections. Their fully convolutional structure enables stable training and parallel computation, making them a widely used baseline for time-series forecasting. In our experiments, the TCN operates directly on the raw multivariate input sequence and produces patch-aligned forecasts under a comparable training protocol.
@@
PatchTST [11] is a Transformer-based forecasting model that applies self-attention to fixed-length temporal patches rather than individual time steps. By treating patches as tokens, PatchTST reduces sequence length and focuses attention on patch-level temporal structure. It serves as a strong baseline for evaluating patch-based Transformer approaches in time-series forecasting.
@@ III-C Implementation details
The proposed model and other compared models have been implemented in python with the library; architecture and experimentation is publicly available [10]. All models are trained using the AdamW optimizer.
@@ III-D Hyperparameter Settings
Tables II and III summarize the global and model-specific hyperparameter settings used in the experiments.
@@
Number of patches ($\lfloor T/P\rfloor$) 20
@@ IV Results
IV-A Synthetic Dataset
All results reported in this section are obtained using a dataset of N=10, [^0] N=10{,}000 samples with sequence length $T=160$.
Figure 2 compares the forecasting performance of the proposed model with the TCN and PatchTST baselines. All models achieve stable performance under the larger dataset size and extended sequence length, indicating that the synthetic dataset remains suitable for evaluating patch-based forecasting approaches at scale. Among the evaluated methods, PatchTST attains the lowest error values, achieving an MSE of 0.055 and an MAE of 0.185. The proposed model achieves an MSE of 0.074 and an MAE of 0.194, while the TCN baseline records an MSE of 0.199 and an MAE of 0.306.
@@ IV-A Synthetic Dataset
The observed performance differences can be interpreted in light of the architectural characteristics of the models. PatchTST directly applies self-attention to patch-level representations that are optimized end-to-end for the forecasting objective, which appears advantageous in the considered synthetic setting, particularly under longer input sequences. This tight integration of patch embedding and attention-based modelling enables efficient capture of nonlinear feature interactions.
The TCN baseline exhibits higher error values under the extended sequence length, suggesting limitations in capturing longer-range dependencies when relying solely on convolutional receptive fields. While TCNs effectively model localized temporal structure, their performance degrades relative to attention-based approaches as the temporal context increases.
@@
The proposed two-stage framework demonstrates consistently improved performance over the convolutional baseline and remains competitive with the patch-based Transformer model. By decoupling local temporal representation learning from global dependency modelling, the approach produces structured patch-level tokens that remain informative for downstream forecasting, even though representation learning and forecasting are performed in separate stages. These results indicate that structured patch-level tokenization continues to be effective as sequence length and dataset scale increase.
IV-B Real-World Evaluation: Electricity Load Forecasting
To assess the generalization capability of the proposed framework beyond controlled synthetic settings, we further evaluate the model on the Electricity Load Diagrams 2011–2014 dataset from the UCI Machine Learning Repository [14]. This dataset consists of electricity consumption measurements recorded at 15-minute intervals for a large number of individual customers over multiple years and is widely used as a benchmark for long-horizon time-series forecasting.
Following standard practice, the data are normalized on a per-meter basis and segmented into sliding windows of fixed length. Each input sample consists of a multivariate sequence of length $T=160$, with the forecasting target defined as the patch-level average load of a selected meter. The same patch size and forecasting horizon used in the synthetic experiments are retained, enabling a consistent evaluation protocol across datasets. Importantly, the convolutional patch encoder trained on synthetic data is reused without architectural modification, highlighting the length-agnostic and transferable nature of the learned local temporal representations.
The proposed model achieves a Mean Squared Error (MSE) of 0.0136 and a Mean Absolute Error (MAE) of 0.0624 on the test split of the Electricity dataset. These results indicate accurate short-horizon forecasting under a real-world setting characterized by strong periodicity, noise, and long-range temporal dependencies. The low MAE value, measured on normalized data, suggests that the model effectively captures both short-term load fluctuations and broader consumption patterns.
The strong performance on electricity load data can be attributed to the architectural separation between local and global modelling stages. Electricity consumption exhibits smooth intra-day dynamics combined with repeating daily and weekly cycles. The convolutional patch encoder operates on short temporal segments, suppressing local noise and extracting stable short-range patterns, while the Transformer focuses on modelling dependencies across patches at a coarser temporal resolution. This division of labour allows the attention mechanism to operate on a more structured and compact representation, which appears particularly effective when longer temporal contexts are considered.
These results demonstrate that the proposed patch-based tokenization strategy generalizes beyond synthetic data and remains effective on a large-scale real-world benchmark. While a detailed comparison with other state-of-the-art electricity forecasting models is beyond the scope of this study, the observed performance provides empirical evidence that decoupling local temporal representation learning from global attention-based modelling yields robust and scalable forecasting behaviour in practical time-series applications.
@@ V Conclusion
This study presents a multi-stage forecasting architecture that explicitly separates patch-level temporal tokenization from global sequence modelling in multivariate time-series forecasting. By learning compact patch-level token embeddings through a convolutional encoder and refining them using token-level self-attention prior to Transformer-based modelling, the proposed framework captures localized temporal structure while enabling effective modelling of longer-range dependencies at a coarser temporal resolution.
Empirical results on a large synthetic dataset with extended sequence length show that the proposed approach consistently outperforms a purely convolutional baseline and achieves competitive performance relative to a strong patch-based Transformer model. Additional evaluation on a real-world electricity load forecasting benchmark demonstrates that the learned patch-level representations transfer effectively beyond controlled synthetic data. While the proposed method does not outperform PatchTST in all settings, it exhibits improved robustness over convolutional baselines and stable performance under increased temporal context. A limitation of this study is the reliance on synthetic data for primary analysis and the use of separate training stages for representation learning and forecasting. Future work will explore jointly trained architectures, alternative patch-level encoders, and broader real-world benchmarks to further assess the generality of the proposed framework.
V-A Limitations and Future Work
A limitation of this study is that a substantial portion of the analysis relies on synthetic data, which, while enabling controlled investigation of model behaviour, may not fully capture the complexity and regime variability present in real-world time-series. Although additional evaluation on electricity load data demonstrates promising generalization, further validation on diverse real-world benchmarks is required.
In addition, the convolutional patch encoder and Transformer components are trained in separate stages, which may restrict the extent to which task-specific information is incorporated into the learned representations. Future work will explore jointly trained architectures, alternative patch-level encoders, and adaptive tokenization strategies to further improve representation quality and forecasting performance.